# Double Robust Variance Estimation

### v1

@@ 2.3 Variance Estimation
Variance estimation for doubly robust estimators (1), (2), and (3) is not straightforward, because the variance of each estimator depends on estimating the parameters of the nuisance models. Here, we consider two methods for variance estimation: the commonly used influence function based variance estimator and the empirical sandwich variance estimator. Both variance estimators can be used to construct Wald-typed CIs for the $ACE$.
@@ 2.3.1 Influence function based variance estimator
Estimators (1), (2), and (3) are suggested by the efficient influence function (Luque-Fernandez et al.,, 2018; Gabriel et al.,, 2023). The commonly used influence function based variance estimator has the form (Glynn and Quinn,, 2010; Gruber and Van Der Laan,, 2012)
$\hat{V}(\widehat{DR})=n^{-2}\sum_{i=1}^{n}\hat{I}_{i}^{2}$ (4)
where $\hat{I}_{i}=\hat{e}_{i}^{-1}X_{i}Y_{i}-(1-\hat{e}_{i})^{-1}(1-X_{i})Y_{i}-\{\hat{e}_{i}^{-1}(1-\hat{e}_{i})^{-1}(X_{i}-\hat{e}_{i})\}\{(1-\hat{e}_{i})\hat{Y}_{i}^{1}+\hat{e}_{i}\hat{Y}_{i}^{0}\}-\widehat{DR}$. Here, pseudo-outcomes $\hat{Y}_{i}^{x}$ equal $a_{x}(Z_{i},\hat{\gamma})$, $b_{x}(Z_{i},\hat{\beta})$, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for the classic AIPW, weighted regression AIPW, and TMLE methods, respectively, and $\widehat{DR}$ equals (1), (2), and (3) for the three corresponding estimators.
The estimators (4) are consistent for the variance of (1), (2), and (3) when both propensity score and outcome models are correctly specified (Daniel,, 2014; Gruber and Van Der Laan,, 2012). When one of the two nuisance models is misspecified, the influence function based variance estimator is not consistent. As noted in the introduction, the estimator (4) is conservative when the weight model is correctly specified but the outcome model is misspecified (Muñoz and Van Der Laan,, 2012), and may be conservative or anti-conservative when the outcome model is correctly specified but the weight model is misspecified (Laan and Robins,, 2003; Muñoz and Van Der Laan,, 2012). In empirical studies, the variance estimator (4) for the classic AIPW estimator (1) was found to be unstable for small sample sizes ($n<1000$), even under correct model specification (Funk et al.,, 2011). Similar instability has been reported for (4) with TMLE ($n=100$) (van der Laan and Gruber,, 2011).
@@ 2.3.2 Empirical sandwich variance estimator
An alternative to the influence function based variance estimator (4) is the empirical sandwich variance estimator. The estimators (1), (2), and (3) can each be expressed as the solution to a set of unbiased estimating equations. The M-estimator for each set of estimating equations, $\hat{\theta}$, is the solution (for $\theta$) to $\sum_{i=1}^{n}\psi_{q}(O_{i},\theta)=0$, where $\psi_{q}(O_{i},\theta)$ are the estimating functions corresponding to each of the three estimators $q\in\{1,2,3\}$ with parameter vector $\theta$. Parameter vectors and estimating functions for (1), (2), and (3) can be specified as follows.
@@
where $\psi_{\beta}$ is the vector of IPTW-weighted score functions for the outcome model. The estimating functions $\psi_{\mu_{WR1}}$ and $\psi_{\mu_{WR0}}$ are for the causal means $\mu_{WR}^{1}$ and $\mu_{WR}^{0}$, respectively, and $\psi_{WR}$ is a delta method transformation of the causal means for estimation of the $ACE$ using the weighted AIPW estimator $\widehat{DR}_{WR}$. Gabriel et al., (2023) demonstrate unbiasedness of this set of estimating equations when at least one of the nuisance models is correctly specified.
@@
With TMLE, $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$. For continuous outcomes, the estimating function $\psi_{3}$ is specified as

### v2

@@ 2.3 Variance Estimation
Variance estimation for doubly robust estimators (1), (2), and (3) is not straightforward, because the variance of each estimator depends on estimating the parameters of the working models. Here, three methods for variance estimation are considered, each of which can be used to construct Wald-typed CIs for the $ACE$.
@@ 2.3.1 Influence function based variance estimator
Estimators (1), (2), and (3) are suggested by the efficient influence function (Luque-Fernandez et al.,, 2018; Gabriel et al.,, 2023). The commonly used influence function based variance estimator has the form $\hat{V}(\widehat{DR})_{IF}=n^{-2}\sum_{i=1}^{n}\hat{I}_{i}^{2}$, where $\hat{I}_{i}=\hat{e}_{i}^{-1}X_{i}Y_{i}-(1-\hat{e}_{i})^{-1}(1-X_{i})Y_{i}-\{\hat{e}_{i}^{-1}(1-\hat{e}_{i})^{-1}(X_{i}-\hat{e}_{i})\}\{(1-\hat{e}_{i})\hat{Y}_{i}^{1}+\hat{e}_{i}\hat{Y}_{i}^{0}\}-\widehat{DR}$ (Gruber and Van Der Laan,, 2012). Here, pseudo-outcomes $\hat{Y}_{i}^{x}$ equal $a_{x}(Z_{i},\hat{\gamma})$, $b_{x}(Z_{i},\hat{\beta})$, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for the classic AIPW, weighted regression AIPW, and TMLE methods, respectively, and $\widehat{DR}$ equals (1), (2), and (3) for the three corresponding estimators.
The influence function based variance estimators $\hat{V}(\widehat{DR})_{IF}$ are consistent for the variance of (1), (2), and (3) when both propensity score and outcome models are correctly specified (Daniel,, 2014; Gruber and Van Der Laan,, 2012). When one of the two working models is misspecified, the influence function based variance estimator is not consistent. As noted in the introduction, the estimator $\hat{V}(\widehat{DR})_{IF}$ is conservative when the weight model is correctly specified but the outcome model is misspecified, and may be conservative or anti-conservative when the outcome model is correctly specified but the weight model is misspecified.
@@ 2.3.2 Empirical sandwich variance estimator
The estimators (1), (2), and (3) can each be expressed as the solution to a set of unbiased estimating equations. The M-estimator for each set of estimating equations, $\hat{\theta}$, is the solution (for $\theta$) to $\sum_{i=1}^{n}\psi_{q}(O_{i},\theta)=0$, where $\psi_{q}(O_{i},\theta)$ are the estimating functions corresponding to each of the three estimators $q\in\{1,2,3\}$ with parameter vector $\theta$. Parameter vectors and estimating functions for each estimator are provided in Appendix A.
[…]
2.3.3 Nonparametric Bootstrap
The nonparametric bootstrap can also be used to estimate the variance of (1), (2), and (3). While variations exist, the nonparametric bootstrap typically involves selecting $B$ independent, with replacement resamples of size $n$ from the observed sample. The estimator of interest is applied to each resample $b$ to obtain $\widehat{DR}_{b}$, where $\widehat{DR}_{b}$ equals (1), (2), or (3) applied to resample $b$. Then, $\hat{V}(\widehat{DR})_{NB}=\sum_{b=1}^{B}\{\widehat{DR}_{b}-\widehat{DR}^{*}\}^{2}/(B-1)$, where $\widehat{DR}^{*}=B^{-1}\sum_{b=1}^{B}\widehat{DR}_{b}$. For large samples, the nonparametric bootstrap is expected to provide valid estimates of the variance for smooth functions of solutions to smooth estimating equations (Davison,, 1997), including (1), (2), and (3). Therefore, the nonparametric bootstrap is a doubly robust variance estimator. However, the bootstrap’s performance may suffer in certain settings, e.g., in the presence of outliers (Davison,, 1997).
3 Example: Improving Pregnancy Outcomes with Progesterone
The doubly robust estimators described in Section 2.2 were applied to IPOP data for estimation of the effect of maternal anemia on birth weight. The exposure, maternal anemia, was defined as having a baseline hemoglobin level below 10.5 g/dL. The IPOP data were limited to the 782 participants (98%) with no prior preterm births and with measured birth weights and baseline hemoglobin values. Note 14 participants experienced stillbirth. Birth weights for these participants were included in the analysis; we return to this issue below. It was assumed that potential outcomes are independent of the exposure conditional on the covariates listed in Appendix Table A1.
To demonstrate differences in the variance estimators presented in Section 2.3, each doubly robust estimator was applied with three sets of model specifications. First, the covariates in Appendix Table A1 were included in both propensity and outcome models. Then, a naive outcome model was fit that included only the exposure variable, while the propensity model included the full set of covariates. Under this specification, the outcome model was reasonably thought to be misspecified due to its simplicity. Finally, a naive propensity model was fit that included only an intercept term, i.e., assuming the propensity score is constant, while the outcome model included the full set of covariates. This propensity model was similarly thought to be misspecified. For all estimators, maternal height was modeled using restricted cubic splines with four knots placed at the 5th, 35th, 65th, and 95th percentiles. The empirical sandwich variance estimator was computed using geex in R (Saul and Hudgens,, 2020) and delicatessen in Python (Zivich et al., 022c, ). The bootstrap was based on 5000 resamples. Point estimates and corresponding 95% CIs for each method and model specification are presented in Table 1, with 95% CI half-widths compared in Figure 1.
The doubly robust methods provided similar point estimates of the $ACE$ for all model specifications, with estimates of approximately -40g, though all 95% CIs included zero. Precision estimates based on the empirical sandwich variance estimator and the influence function based variance estimator were similar when the full covariate set was included in both working models. Under the naive outcome model, standard errors for the influence function based variance estimator were larger than those for the empirical sandwich variance estimator. When the propensity model was naive, standard errors for the influence function based variance estimator were smaller than those for the empirical sandwich variance estimator. Note there was more fluctuation in 95% CI half-widths across model specifications for the influence function based variance estimator than the empirical sandwich variance estimator (Figure 1). Bootstrap standard errors were larger than those of the empirical sandwich and influence function based variance estimators. This may be due to the presence of extreme birth weights in the data (Appendix Figure A1).
There are notable limitations associated with this analysis. First, the $ACE$ is likely not the most informative estimand to address this research question. The $ACE$ contrasts average birth weight under settings where all women have anemia and no women have anemia. It is challenging to envision a setting where all women would be susceptible to anemia, so the estimators (1), (2), and (3) may lack a meaningful causal interpretation. While these statistical quantities might still be beneficial, more informative estimands could also be considered to address this research question, including a comparison of $E(Y^{0})$ with the natural course, $E(Y)$ (Hubbard and Van der Laan,, 2008) or contrasts of stochastic intervention distributions (Kennedy,, 2019). Additionally, as with all observational studies, there may be uncontrolled confounding not captured in the observed set of covariates. This concern may be heightened in settings where causal consistency is also questionable (Hernán and VanderWeele,, 2011), i.e., there may be multiple interventions to modify anemia status. Finally, the analysis included participants who experienced stillbirth. Future work could reexamine this problem using alternative approaches for competing events.
$\widehat{DR}$ ES-SE ES-95% CI IF-SE IF-95% CI NB-SE NB-95% CI
Full Covariate Set
Classic AIPW -37 56 (-147, 73) 58 (-151, 76) 63 (-161, 86)
Wtd regression AIPW -41 56 (-151, 69) 56 (-151, 68) 61 (-160, 78)
TMLE -37 56 (-147, 73) 58 (-151, 76) 63 (-160, 85)
Naive Outcome
Classic AIPW -36 57 (-148, 76) 61 (-156, 84) 64 (-162, 90)
Wtd regression AIPW -36 57 (-148, 77) 61 (-156, 84) 64 (-161, 90)
TMLE -36 57 (-148, 77) 61 (-156, 84) 63 (-160, 89)
Naive Propensity
Classic AIPW -41 58 (-154, 73) 56 (-150, 69) 58 (-155, 74)
Wtd regression AIPW -47 57 (-159, 65) 54 (-153, 59) 59 (-162, 68)
TMLE -41 58 (-154, 73) 56 (-150, 69) 59 (-155, 74)
  • Note: SE=standard error, ES=empirical sandwich, IF=influence function, NB=nonparametric bootstrap, CI=confidence interval, Wtd=weighted. The naive outcome model included only the exposure, and the naive propensity model included only an intercept. The nonparametric bootstrap was based on 5000 resamples.
4 Simulation Study
A simulation study was conducted to compare the empirical properties of the influence function based variance estimator, the empirical sandwich variance estimator, and the nonparametric bootstrap in conjunction with each of the three doubly robust methods discussed in Section 2. Simulations were conducted with $n=800$, similar to the sample size in the example in Section 3.
4.1 Simulation Setup
Three covariates ($Z_{1}$, $Z_{2}$, $Z_{3}$), the exposure ($X$), and potential outcomes ($Y^{0}$ and $Y^{1}$) were simulated. The covariate $Z_{1}$ was distributed Normal with mean $155$ and standard deviation $7.6$. Two binary covariates $Z_{2}$ and $Z_{3}$ were simulated from Bernoulli distributions with means $0.25$ and $0.75$, respectively. The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(15-0.1Z_{1}+2.5Z_{2}-Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$. Potential outcomes $Y^{1}$ and $Y^{0}$ under exposure and no exposure, respectively, were simulated from a normal distribution with mean $E(Y^{x})=1000+11.5Z_{1}+100Z_{2}-15Z_{1}Z_{2}+25x-5.5xZ_{1}-30xZ_{2}+5xZ_{1}Z_{2}$ and standard deviation $\sigma=400$ for $x\in\{0,1\}$. Under this data generating mechanism, the $ACE$ was approximately $-60$. The marginal distributions of the exposure and outcome were modeled after the example in Section 3. To examine performance of the estimators under the null, simulations were also conducted where $E(Y^{1})$ was equal to $E(Y^{0})$, as defined above, such that $ACE=0$.
The estimators $\widehat{DR}_{C}$, $\widehat{DR}_{WR}$, and $\widehat{DR}_{TMLE}$ were applied to 5000 simulated samples. Estimators were computed four ways: (1) propensity and outcome models correctly specified, (2) propensity models correctly specified but outcome models misspecied, (3) outcome models correctly specified but propensity models misspecified, and (4) both models misspecified. Misspecified propensity models included only an intercept and a linear term for $(Z_{1}-155)^{2}$, and misspecified outcome models included only an intercept and linear terms for $X$ and $(Z_{1}-155)^{2}$.
Each of the three variance estimators described in Section 2.3 was applied. For the nonparametric bootstrap, 1000 resamples were included in each iteration for estimation of the bootstrap standard error, excluding any resamples where working models failed to converge. Simulation results were summarized by empirical bias, average standard error (ASE), empirical standard error (ESE), standard error ratio (SER=ASE/ESE), and empirical 95% CI coverage. The ratio of the variance estimate for each simulation and the ESE, i.e., the variance ratio, was also summarized. That is, $VR=se_{s}/ESE$, where $se_{s}=\sqrt{\hat{V}(\widehat{DR})}$ for a given doubly robust estimator, model specification, and variance estimator for simulation $s$.
4.2 Simulation Results
Both point and variance estimators performed as expected in simulations. The results of the simulation study are presented in Table 2 and Figures 2 and 3. When at least one model was correctly specified, the classic AIPW, weighted regression AIPW, and TMLE displayed minimal bias, but all were substantially biased when both working models were misspecified. Under correct specification of both working models, all three variance estimators tracked closely with the ESE, resulting in SERs close to one and CIs attaining the nominal level of coverage. When both models were misspecified, all three variance estimators tracked with the ESE, but bias was substantial, resulting in CIs with below nominal coverage.
Differences between variance estimators are apparent for scenarios where only one working model was correctly specified. When either the outcome model or the propensity model was misspecified, CIs based on the empirical sandwich variance estimator and the nonparametric bootstrap attained the nominal level of coverage, and estimated standard errors generally tracked closely with the ESE. In Figures 2 and 3, note variance ratios for the empirical sandwich estimator and the nonparametric bootstrap are clustered around one for these scenarios. In contrast, the influence function based variance estimator was empirically biased when either working model was misspecified. As expected, it tended to overestimate the variance when the outcome model was misspecified, leading to SERs above one and conservative CI coverage. When the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator underestimated the variance, resulting in SERs below one. Bias in the influence function based variance estimator was more pronounced under the null. Under outcome model misspecification, SERs exceeded 2.0, resulting in approximately 100% CI coverage. Under propensity misspecification, SERs were 0.89, resulting in CIs with below nominal coverage.
The empirical sandwich variance and nonparametric bootstrap estimators both demonstrated the expected doubly robust variance property, but there was a notable difference in the performance of the estimators. The nonparametric bootstrap generally demonstrated more variation than the empirical sandwich estimator, as evidenced by increased spread and more extreme outliers in Figures 2 and 3.
Additional simulations were conducted with $\sigma\in\{200,600\}$, with a binary outcome, and with $n=2000$. Twelve scenarios were considered. Details and results of the additional scenarios are provided in Appendix C.
In summary, the simulation study demonstrated the theoretical properties explained in Section 2. That is, the empirical sandwich variance estimator and the nonparametric bootstrap were empirically unbiased when at least one of the two working models was correctly specified. The influence function based variance estimator is not consistent under misspecification of either working model. The influence function based variance estimator was conservative when the outcome model was misspecified and was generally anti-conservative when the propensity model was misspecified. The magnitude and direction of bias under misspecified working models varied across simulation scenarios.
Figure 2: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=800, $\sigma=400$, 5000 simulations. $ACE$ was approximately -60. Black squares denote the mean variance ratio (=SER). Results exclude one simulation where models failed to converge. The $0.33\%$ of correct model specification simulations, $4.02\%$ of misspecified outcome model simulations, and $0.004\%$ of misspecified propensity model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
Figure 3: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=800, $\sigma=400$, 5000 simulations under the null. Black squares denote the mean variance ratio (=SER). The $0.01\%$ of correct model specification simulations and $1.72\%$ of misspecified outcome model simulations where the ratio was above 2.75 or below 0.5 are not displayed.
Table 2: Simulation summary results, continuous outcome, $n=800$, $\sigma=400$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
$ACE\approx-60$
CS Classic 0.4 58.4 0.99 95 1.00 95 1.00 95
WR 0.4 58.3 0.99 95 1.00 95 0.99 95
TMLE 0.4 58.3 0.99 95 1.00 95 1.00 95
MO Classic -0.4 60.0 0.99 95 1.02 96 1.07 97
WR -1.6 59.2 0.99 95 1.00 95 1.05 96
TMLE -0.5 59.7 0.99 95 1.01 95 1.06 96
MP Classic 0.3 57.8 1.00 95 1.00 95 0.97 94
WR 0.3 57.8 1.00 95 1.00 95 0.97 94
TMLE 0.3 57.8 1.00 95 1.00 95 0.97 94
MB Classic -23.8 57.0 1.00 92 1.00 92 1.00 92
WR -23.8 57.0 1.00 92 1.00 92 1.00 92
TMLE -23.8 57.0 1.00 92 1.00 92 1.00 92
$ACE=0$
CS Classic 0.1 35.0 0.96 94 1.00 95 0.97 94
WR 0.1 34.8 0.96 94 0.98 94 0.95 94
TMLE 0.1 35.0 0.96 94 0.99 95 0.97 94
MO Classic 1.4 48.0 0.93 94 1.07 96 2.10 100
WR 2.5 44.8 0.94 93 1.02 95 2.18 100
TMLE 1.6 46.7 0.93 94 1.04 96 2.12 100
MP Classic 0.0 33.7 0.98 95 0.99 95 0.89 92
WR -0.1 33.7 0.98 95 0.99 95 0.89 92
TMLE 0.0 33.7 0.98 95 0.99 95 0.89 92
MB Classic 158.8 75.9 0.99 44 0.99 44 0.99 44
WR 158.8 75.9 0.99 44 0.99 44 0.99 44
TMLE 158.8 75.9 0.99 44 0.99 44 0.99 44
  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage, i.e., the proportion of simulated samples for which the 95% CI included ACE; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%. Results exclude 1 simulation where models did not converge.
5 Discussion
Doubly robust estimators have gained popularity due to their ability to provide consistent point estimates when either an outcome or propensity model is correctly specified. Here, the commonly-used influence function based variance estimator is compared with the empirical sandwich variance estimator and the nonparametric bootstrap in conjunction with three doubly robust estimators: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE. For estimation of the average causal effect with observational data, the influence function based variance estimator is consistent only when both outcome and propensity models are correctly specified. In contrast, both the empirical sandwich variance estimator and the nonparametric bootstrap are doubly robust variance estimators. As such, confidence intervals constructed from these estimators are expected to provide nominal CI coverage when either model is correctly specified.
This paper only considers variance estimation of the average causal effect with observational data. The influence function based variance estimator can be consistent under misspecification of one working model in some settings. For example, this variance estimator is consistent under outcome model misspecification in the randomized setting under additional assumptions (Williams et al.,, 2022; Van Lancker et al.,, 2024) and for estimation of some non-degenerate estimands (Haneuse and Rotnitzky,, 2013).
While here consideration was limited to finite-dimensional parametric modeling approaches, machine learning approaches are commonly applied with doubly robust estimators. These approaches allow for more complex functional forms for continuous covariates than fully parametric approaches and include higher order interaction terms that may not be included in investigator-specified parametric models, often leading to estimators that are more robust to model misspecification (Zivich et al.,, 2022). However, convergence of machine learning algorithms is typically slower and consistent variance estimation is more challenging than with parametric modeling approaches. Machine learning methods are typically not compatible with the estimating equations approach discussed in this paper, though alternative methods have been developed for doubly robust variance estimation in this context for TMLE estimators (Benkeser et al.,, 2017).
Use of the empirical sandwich variance estimator in conjunction with doubly robust estimators is not new, though its doubly robust property has not been emphasized. Lunceford and Davidian, (2004) discuss use of the empirical sandwich variance estimator for estimating the variance of the classic AIPW estimator and note that it tends to be more stable than the influence function based variance estimator. However, based on the systematic review discussed in the introduction (Smith et al.,, 2023), the empirical sandwich variance estimator does not appear to be widely used with TMLE. While some existing software packages compute empirical sandwich variance estimators for doubly robust estimators (e.g., the “causaltrt” procedure in SAS and the “dr” procedure in Stata), other popular software packages compute the influence function based variance estimator but not the empirical sandwich variance estimator (e.g., the AIPW and tmle packages in R and the zEpid package in Python). We hope this work will allow for easier implementation of doubly robust point and variance estimation.
Supplementary Material
Acknowledgements
This work was supported by the National Institutes of Health under award numbers R01 AI157758, K01 AI182506, R01 AI085073, P30 AI050410, and K01 AI177102. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH. The authors thank Drs. Bradley Saul and Michael Hudgens at the University of North Carolina for coding support and helpful suggestions, respectively.
Appendix A: parameter vectors and estimating functions for three doubly robust estimators with the M-estimation approach
@@ Acknowledgements
where $\psi_{\beta}$ is the vector of IPTW-weighted score functions for the outcome model and $b_{x}(Z_{i},\beta)=\phi^{-1}(H_{i}(x,Z_{i}))^{\top}\beta$ for $x\in\{0,1\}$. The estimating functions $\psi_{\mu_{WR1}}$ and $\psi_{\mu_{WR0}}$ are for the causal means $\mu_{WR}^{1}$ and $\mu_{WR}^{0}$, respectively, and $\psi_{WR}$ is a delta method transformation of the causal means for estimation of the $ACE$ using the weighted regression AIPW estimator $\widehat{DR}_{WR}$. Gabriel et al., (2023) demonstrate unbiasedness of this set of estimating equations when at least one of the working models is correctly specified.
@@
For TMLE, $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$. For continuous outcomes, the estimating function $\psi_{3}$ is specified as